# Learning Koopman Dynamics for Safe Legged Locomotion with Reinforcement Learning-based Controller

### v1

@@ IV-B Data collection for high-level dynamics
We create the dataset of high-level floating base movements by generating 100 locomotion trajectories. Each trajectory consists of 10 seconds of robot state recorded at 50 Hz, which results in 500 data points for each trajectory. We obtain diverse and dynamic trajectories by randomly sampling velocity commands every 0.5 seconds. For each trajectory, we record the high level states of the robot state: $\mathrm{x}=[p_{x},p_{y},\theta,v_{x},v_{y},\omega]$, where $p_{x},p_{y},\theta$ are $x,y$ coordinates and direction heading of the robot and $v_{x},v_{y},\omega$ are $x,y$ linear velocity components and $z$-axis angular velocity of the robot. We also record the velocity tracking commands given to the robot: $\mathrm{u}=[\hat{v_{x}},\hat{v_{y}},\hat{\omega}]$, where the commands are relative to the position and orientation of robot’s state at each step.
For better coverage, we apply a data augmentation technique by applying transformations to the existing dataset. For each episode of length $N$ state transitions $[\mathrm{x}_{0},\mathrm{u}_{0},\dots,\mathrm{u}_{N-1},\mathrm{x}_{N}]$, we extract sequences of length $H$, leading to $(N-H+1)$ sequences per $N$ step trajectory. For each sequence of length $H$, we convert the positional and angular components to the coordinate of its first state.
Formally, for $i$-th state sequence of an episode: $[\mathrm{x}_{i},\dots,\mathrm{x}_{i+H}]$, we convert each state represented in a global coordinate to the local coordinate system with respect to the first state $\mathrm{x}_{i}$. This will lead to a representation of $[\mathrm{x}^{i}_{i},\dots,\mathrm{x}^{i}_{i+H}]$, where $\mathrm{x}^{i}_{j}$ represents a $\mathrm{x}_{j}$ represented in the frame of $\mathrm{x}_{i}$. Similarly for the command sequence $[\mathrm{u}_{i},\dots,\mathrm{u}_{i+H-1}]$, we convert each velocity commands in each step’s robot coordinate to the coordinate of the first step: $[\mathrm{u}^{i}_{i},\dots,\mathrm{u}^{i}_{i+H-1}]$. As a result, we obtain the final dataset $\mathcal{D}$ which contains $2,255,000$ frames of state transitions.
@@ IV-C Learning high-level dynamics
One key challenge in learning $F$ is that it can be arbitrarily nonlinear. One potential way in machine learning to learn the transition function $F$ is to utilize a universal function approximator, a multi-layer perceptron (MLP), to model the input-output mapping of the function $F$ minimizing the $L_{2}$ loss: $||x_{t+1}-\text{MLP}(x_{t},u_{t})||_{2}$ for $\forall t$. However, utilizing MLP dynamics constraint in the MPC framework causes practical challenges, which makes MLP-learned dynamics impractical for general usage. For instance, Direct Multiple Shooting formulation [45, 46, 28], one of the widely used techniques in solving MPC, requires the dynamics constraint containing MLP to match for every time step, making the optimization extremely nonlinear. As a result, nonlinear solvers often fail to solve the optimization with MLP dynamics constraints.
Instead, we use the Koopman operator’s linear evolution to learn the linear evolution in the lifted space, which makes the dynamics constraint in the MPC framework linear. As a result, our approach offers the possibility of adopting linear MPC with control barrier functions. In this work, we use the Koopman forward dynamics as described in Section. III: $\phi(\mathrm{x}_{t+1})=A\phi(\mathrm{x}_{t})+Bu_{t}$, where the state-lifting, $\phi(\cdot):R^{n}\to R^{n^{\prime}}$ $n^{\prime}\geq n$, is a vector-valued lifting function. Inspired by [33], the first $n$ elements of the $\phi(\mathrm{x}_{t})$ are $\mathrm{x}_{t}$ so that we can easily retrieve the original states, which are used for computing obstacle-avoidance constraints. Note that in the case of $n^{\prime}=n$, $\phi(\mathrm{x})$ becomes the same as $\mathrm{x}$, and the solutions $A$ and $B$ becomes linear identification of the high-level dynamics. Finally, we obtain the analytical solution $A$ and $B$ matrices by solving the optimal solution of $J(K)$ in Equation. 2, which in general only takes seconds of cpu-only computation.
IV-D Obstacle avoidance with Koopman MPC
Once we obtain the learned linear evolution in the lifting space, we can incorporate the existing control framework for high-level tasks. In our case, we leverage the existing linear Model-Predictive Control algorithm to design the safety-guaranteed legged navigation controller. Our optimization is formulated as follows:
@@ IV-D Obstacle avoidance with Koopman MPC
The key idea of our framework is to directly solve lifted variables $\phi_{1\cdots H}$, instead of the original states $\mathrm{x}_{1\cdots H}$, along with control variables $\mathrm{u}_{1\cdots H}$. This formulation is effective because all the dynamics become linear in the lifted space. Therefore, we can easily solve the given optimization using Quadratic Programming.
Our objective function is design to track the given trajectory $\hat{x}_{1\cdots H}$ in the original state spaces:
$J(\mathrm{\phi}_{1\cdots H},\mathrm{u}_{1\cdots H})=\sum_{k=1}^{H}||\mathrm{x_{k}}-\mathrm{\hat{x}_{k}}||_{P}+\sum_{k=0}^{H-1}||\mathrm{u}_{k+1}-\mathrm{u}_{k}||_{2},$
where the $\mathrm{x_{k}}=\mathrm{\phi^{-1}}(\phi_{k})$ for all the timestep $k$. Because the lift function $\phi$ to include the original state information, the inverse operation $\phi^{-1}(\cdot)$ is a linear projection, making the optimization to be quadratic with respect to all the variables.
The first constraint aims to satisfy the learned dynamics. The second constraint is an arbitrary constraint, in our case, an distance barrier function $h_{i}(\phi_{k})=\hat{h}_{i}(\mathrm{x}_{k})=\textbf{dist}(\mathrm{x}_{k},\text{obstacle}_{i})$ for all obstacles. The third constraint confines control signals within the valid region.
@@ V Experimental Evaluation
We evaluated the our method along with baselines in terms of the long-term forward prediction accuracy (Sec. V-B) and the MPC-based safe navigation performance (Sec. V-C).
@@ V-A Experimental Design
Evaluation Platform: We collected all locomotion data and tested the navigation performance using the Unitree’s Aliengo [48] robot built with IsaacGym simulator [49].
Dynamics Models: We compared various Koopman Dynamics models along with the baselines:
Linear Integrator: An ideal first order linear forward dynamical model: $[p_{x},p_{y},\theta]_{t+1}=[p_{x},p_{y},\theta]_{t}+\frac{dt}{2}[v_{x},v_{y},\omega]_{t}+\frac{dt}{2}[\hat{v_{x}},\hat{v_{y}},\hat{\omega}]_{t}$; $[v_{x},v_{y},\omega]_{t+1}=[\hat{v_{x}},\hat{v_{y}},\hat{\omega}]_{t}$, where $dt=0.02s$ represents the duration of one simulation step. Note that this baseline does not require data for training.
Component-wise Linear Dynamics: Originally proposed in [19], this baseline assumes a linear forward dynamics model for each state component, such as $[p_{x},v_{x}]$ and $\hat{v_{x}}$. Therefore, it has three pairs of matrices $[\bar{A}\in\mathbb{R}^{2\times 2},\bar{B}\in\mathbb{R}^{2\times 1}]$, which are all learned from the dataset $\mathcal{D}$.
Unified Linear Dynamics: This baseline assumes $\phi(\mathrm{x})=\mathrm{x}$ (i.e., no other lifting functions), indicating a linear system evolution under control inputs in the original state space. Similarly, these matrices are learned from the dataset $\mathcal{D}$.
Koopman (P): Inspired by [33], the Koopman model used the lifting function $\phi(\mathrm{x}_{t})$ as polynomials of up to degree three.
Koopman (TD(n)): Another commonly used lifting function is Time-Delay Embedding [50], which concatenates the current system state with the previous $n-1$ history states, i.e., $\phi(\mathrm{x}_{t})=[\mathrm{x}_{t-n+1};\cdots;\mathrm{x}_{t}]$. Note that $\mathrm{u}_{t}$ still remains a single step control command, as required by the MPC formulation. We evaluated different $n$ values: 5, 10, 20, and 30.
@@ V-B Long-term Forward Prediction
Integrator Comp. Linear Unified Linear Koopman (P) Koopman (TD(5)) Koopman (TD(10)) Koopman (TD(20)) Koopman (TD(30))
ValidationSet 0.40($\pm$0.39) 0.27($\pm$0.24) 0.07($\pm$0.09) 0.07($\pm$0.09) 0.07($\pm$0.09) 0.06($\pm$0.09) 0.05($\pm$0.08) 0.04($\pm$0.08)
NewSet1 0.41($\pm$0.49) 0.19($\pm$0.19) 0.09($\pm$0.11) 0.09($\pm$0.11) 0.09($\pm$0.11) 0.09($\pm$0.11) 0.07($\pm$0.10) 0.05($\pm$0.09)
NewSet2 0.39($\pm$0.35) 0.29($\pm$0.26) 0.06($\pm$0.08) 0.06($\pm$0.08) 0.06($\pm$0.07) 0.05($\pm$0.07) 0.04($\pm$0.07) 0.03($\pm$0.06)
NewSet3 0.40($\pm$0.33) 0.32($\pm$0.27) 0.05($\pm$0.07) 0.05($\pm$0.07) 0.05($\pm$0.07) 0.05($\pm$0.07) 0.04($\pm$0.06) 0.03($\pm$0.05)
TABLE I: Average Prediction Error for each dataset, calculated across all 10,000 randomly selected sequences, states, and time steps.
We first evaluate the long-term forward prediction accuracy of each model with respect to the Prediction Error, which measures the absolute differences between predicted and ground-truth states. For a fair comparison, all models were trained on the same dataset, except for the Linear Integrator. We then rolled out the same PPO controller with random velocity commands to collect a new set of 100 locomotion trajectories ($T=500$), which were converted into length-$100$ sequences as the validation dataset.
Figure 3: Average Prediction Error over 10,000 selected sequences. Solid lines represent mean values, and shaded areas indicate standard deviation.
In Fig. 3, we present the average Prediction Error over 10,000 randomly selected sequences for each state at each time step. From the results, it is evident that the Koopman (TD(30)) model significantly outperforms the other models in terms of long-term prediction accuracy. Additionally, its prediction errors are nearly zero within the span of the time-delay embedding. However, surprisingly, the Unified Linear Dynamics and Koopman (P) exhibit very similar performance, as the green and purple lines are nearly indistinguishable, as shown in Fig. 3, suggesting that the polynomial liftings may not effectively capture the underlying dynamics.
We then adjusted the velocity sampling interval (default: 0.5s, used for both training and validation) to 0.1s, 1s, and 3s, and collected three additional test datasets, referred to as NewSet1, NewSet2, and NewSet3. In Table. I, we report each model’s average Prediction Error for each dataset, calculated across all 10,000 randomly selected trajectories, states, and time steps. These results show that the Unified Linear Dynamics and Koopman (P) indeed perform very similarly across all datasets. On the other hand, the results also show that the Koopman forward dynamics benefits from larger time-delay embedding size, and can achieve half the prediction error of the Unified Linear Dynamics when $H=30$. This finding aligns with the results reported in [51], where it shows that providing more history states as input to the neural network policy enhances locomotion performance. However, from the standpoint of practical MPC implementation, there is a trade-off between prediction accuracy and computational burden imposed by larger $A$ and $B$ matrices.
@@ V-C MPC-based Safe Navigation
In this section, we utilize the learned Unified Linear Dynamics model in MPC-based safe navigation settings. We omit Koopman (P) or Koopman (TD) models because Unified Linear Dynamics shows reasonable performance (Fig. 3) and is suitable for real-time control. Note that it is straightforward to extend to the Koopman Dynamics models using either function bases or time-delay embeddings as the lifting function because for both we can easily retrieve $\mathrm{x}_{t}$ from $\phi(\mathrm{x})$ using linear projection for MPC objective and safe constraints.
The navigation framework is tested on three different maps with randomized start and goal selection. To decouple the vision and focus on the effectiveness of our planning framework, we assume the positions and shapes of the obstacles are known. High-level planners are designed to find rough waypoints, which are used as reference trajectories for the MPC controller.
Figure 4: Overview of the navigation framework.
@@
Corridor 1 Corridor 2 Maze (width=0.75) Maze (width=0.7)
Time (s) $\downarrow$ Collisions $\downarrow$ Success (%) $\uparrow$ Time (s) $\downarrow$ Collisions $\downarrow$ Success (%) $\uparrow$ Time (s) $\downarrow$ Collisions $\downarrow$ Success (%) $\uparrow$ Time (s) $\downarrow$ Collisions $\downarrow$ Success (%) $\uparrow$
Integrator 5.48 1.65 70 5.65 1.25 55 19.68 4.5 90 20.79 12.5 70
Comp. Linear 4.18 1.95 95 4.26 2 100 15.39 4.8 100 14.99 13.4 100
Unified Linear 4.30 1.3 100 4.31 2.25 95 15.21 3.3 100 14.50 6 100
TABLE II: Safe navigation results. Each value represents the mean over 20 random runs for Corridor environments and 10 random runs for Maze environments.
Navigation Environments: We test the effectiveness of our learned models by deploying the robots in 3 different navigation environments: i) corridor 1: a wider corridor with a 70cm gap, ii) corridor 2: a narrower corridor with a 55cm gap, and iii) maze: an artifact maze [28]. For each of the navigation environments, we randomize start and goal positions. We then randomly sample 10 combination of start position and goal position pairs.
High-Level Planner: To navigate complex environments such as mazes, using a search-based high level planner enables creating rough routes toward goals. For each of the environments, we utilize an A* algorithm [52] to create a global path from the start to the goal. At each step of MPC, we truncate the goal and create a local trajectory based on the current position of the robot.
MPC Controller: Based on the desired local trajectory obtained from the high-level planner, we use MPC to compute velocity commands for the low-level controller. We follow the objective in IV-D. Here, we model the geometry of the robot using three circles with a radius of $25$cm, which are equally distributed along the body. The safety constraint is then computed based on the distance between each circle and the polytope modeling of obstacles. The optimization is solved using IPOPT interfaced through CasADi[53]. It’s worth noting that utilizing linear dynamics learned through multi-layer perceptron leads to poor MPC performance due to nonlinear equality constraints.
@@
Experiment results for Linear Integrator, Component-wise Linear, and Unified Linear models are depicted in Table II. We measure i) seconds taken to reach destination, ii) the number of collisions occurred during navigation, and iii) the success rate of the navigation. Here, the success is determined based on whether the robot was able to reach the destination in a desired time window, where we give 6 seconds for corridors and 30 seconds for maze environments. We did not terminate the episode upon minor collisions. However, when the robot fell down and were not be able to recover, we marked them as failure cases.
Overall, the learned Unified Linear model demonstrates the best performance with the lowest number of collisions. This trend becomes more distinctive when it comes to more challenging environments, such as Maze (width=0.75) and Maze (width=0.7), where the Unified Linear model made collisions near 68% or 44% of the second-best method, the Component-wise Linear model. Therefore, we can conclude that the proposed safe navigation framework shows good performance even in complex environments, which have much more dense obstacles compared to the previous works [27].
However, the Linear Integrator model was often unable to succeed the tasks due to significant collisions, making the body fall down. Or, it faced too much deviation from the desired trajectory that it was not able to find adequate commands. Because we did not record additional collisions after the termination, this model achieved the least collisions in Corridor 2 with the worst success rate of $55$%.
It is worth noting that the Component-wise Linear model performed similarly with the Unified Linear model in the corridor tasks. This was a bit surprising because we initially expect much worse navigation performance based on the model prediction performance analyzed in Sec. V-B. Therefore, we suspect that the estimation performance may not be proportionally transferred to the MPC planning performance, depending on the tasks, particularly when they are relatively easy. This leads to room for exploration toward the validity of weak models on safe navigation tasks.

### v2

@@ IV-B Data Collection for High-level Dynamics
We create the dataset of high-level floating base movements by generating 150 locomotion trajectories. Each trajectory consists of 150 seconds of robot state recorded at 50 Hz, which results in 7500 data points for each trajectory. We obtain diverse and dynamic trajectories by randomly sampling velocity commands every 0.5 seconds. For each trajectory, we record the high level states of the robot state: $\mathrm{x}=[p_{x},p_{y},\theta,v_{x},v_{y},\omega]$, where $p_{x},p_{y},\theta$ are $x,y$ coordinates and direction heading of the robot and $v_{x},v_{y},\omega$ are $x,y$ linear velocity components and $z$-axis angular velocity of the robot. We also record the velocity tracking commands given to the robot: $\mathrm{u}=[\hat{v}_{x},\hat{v}_{y},\hat{\omega}$], where the commands are relative to the position and orientation of robot’s state at each step.
For better coverage, we apply a data augmentation technique by applying transformations to the existing dataset. For each episode of length $N$ state transitions $\mathrm{x}_{0},\mathrm{u}_{0},\dots,\mathrm{u}_{N-1},\mathrm{x}_{N}]$, we extract sequences of length $H$, leading to (NH+1)(N-H+1) sequences per $N$ step trajectory. For each sequence of length $H$, we convert the positional and angular components to the coordinate of its first state.
Formally, for $i$-th state sequence of an episode: $\mathrm{x}_{i},\dots,\mathrm{x}_{i+H}]$, we convert each state represented in a global coordinate to the local coordinate system with respect to the first state $\mathrm{x}_{i}$. This will lead to a representation of $\mathrm{x}^{i}_{i},\dots,\mathrm{x}^{i}_{i+H}]$, where $\mathrm{x}^{i}_{j}$ represents a $\mathrm{x}_{j}$ represented in the frame of $\mathrm{x}_{i}$. As a result, we obtain the final dataset $\mathcal{D}$ which contains $236,250$ frames of state transitions.
@@ IV-C Learning high-level dynamics
One key challenge in learning $F$ is that it can be arbitrarily nonlinear. One potential way in machine learning to learn the transition function $F$ is to utilize a universal function approximator, a multi-layer perceptron (MLP), to model the input-output mapping of the function $F$ minimizing the $L_{2}$ loss: $x_{t+1}-f_{MLP}(x_{t},u_{t})||_{2}$ for $\forall t$. However, utilizing MLP dynamics constraint in the MPC framework causes practical challenges, which makes MLP-learned dynamics impractical for general usage. For instance, Direct Multiple Shooting formulation [42, 43, 25], one of the widely used techniques in solving MPC, requires the dynamics constraint containing MLP to match for every time step, making the optimization extremely nonconvex. As a result, nonlinear solvers often fail to solve the optimization with MLP dynamics constraints.
Instead, we use the Koopman operator’s linear evolution to represent dynamics in the lifted space, which makes the dynamics constraint in the MPC framework linear. As a result, our approach offers the possibility of adopting linear MPC with control barrier functions. In this work, we use the Koopman forward dynamics as described in Section. III: $\phi(\mathrm{x}_{t+1})=A\phi(\mathrm{x}_{t})+Bu_{t}$, where the state-lifting, $\phi(\cdot):R^{n}\to R^{n^{\prime}}$ $n^{\prime}\geq n$, is a vector-valued lifting function. Inspired by [30], the first $n$ elements of the $\phi(\mathrm{x}_{t})$ are $\mathrm{x}_{t}$ so that we can easily retrieve the original states, which are used for computing obstacle-avoidance constraints. Note that in the case of $n^{\prime}=n$, $\phi(\mathrm{x})$ becomes the same as $\mathrm{x}$, and the solutions $A$ and $B$ becomes a linear identification of the high-level dynamics. Finally, we obtain the analytical solution $A$ and $B$ matrices by solving the optimal solution of $J(K)$ in Equation. 1, which in general only takes seconds of CPU-only computation.
IV-D Safe Navigation with Koopman-based MPC
Figure 3: Overview of the navigation framework.
Once we obtain the learned linear evolution in the lifting space, we can incorporate the existing control framework for high-level tasks. In our case, we leverage the existing linear Model-Predictive Control algorithm to design the safety-guaranteed legged navigation controller. The overview is illustrated in Fig. 3. Our optimization can be formulated as follows:
@@ IV-D Safe Navigation with Koopman-based MPC
The key idea of our framework is to directly solve lifted variables $\phi_{1\cdots H}$, instead of the original states $\mathrm{x}_{1\cdots H}$, along with control variables $\mathrm{u}_{1\cdots H}$. This formulation is effective because all the dynamics become linear in the lifted space.
Our objective function is designed to track the given trajectory $\hat{x}_{1\cdots H}$ in the original state space:
[…]
where the $\mathrm{x_{k}}=\mathrm{\phi^{-1}}(\phi_{k})$ for all the timestep $k$. Designing the lifting function $\phi$ to include the original state information, the inverse operation $\phi^{-1}(\cdot)$ becomes a linear projection, making the cost function to be quadratic with respect to all the variables.
The first constraint aims to satisfy the learned dynamics. The second constraint is an arbitrary task-specific constraint, in our case, a distance barrier function $h_{i}(\phi_{k})=\hat{h}_{i}(\mathrm{x}_{k})=\textbf{dist}(\mathrm{x}_{k},\text{obstacle}_{i})$ for all obstacles. The third constraint confines control signals within the valid region.
@@ V Experimental Evaluation
We evaluated our method along with baselines in terms of the long-term forward prediction accuracy (Sec. V-B) and the MPC-based safe navigation performance (Sec. V-C).
@@ V-A Experimental Design
We collected all locomotion data and tested the navigation performance using the Unitree G1 [39] robot built with IsaacGym simulator [41]. The MPC optimization is solved using IPOPT interfaced through CasADi [45].
We compared various Koopman Dynamics models along with the baselines:
Integrator: An ideal first-order model: $p_{x},p_{y},\theta]_{t+1}=[p_{x},p_{y},\theta]_{t}+dt[v_{x},v_{y},\omega]_{t}$; $v_{x},v_{y},\omega]_{t+1}=[\hat{v}_{x},\hat{v}_{y},\hat{\omega}]_{t}$, where $dt=0.02s$ represents the duration of one simulation step and with the velocities accumulated in the global frame. Note that this baseline does not require data for training.
Component-wise Linear (Comp.Lin.): Building on the baseline in [13], we learn a block-structured linear model by assuming independence among the components in the state dynamics.
Unified Linear (Linear): This baseline assumes $\phi(\mathrm{x})=\mathrm{x}$ (i.e., no other lifting functions), indicating a linear system evolution under control inputs in the original state space. Similarly, these matrices are learned from the dataset $\mathcal{D}$.
Model MSE RMSE Training Lifted RMSE RMSE RMSE
Time (s) Dim X (m) Y (m) Yaw (rad)
Koopman Lifting Models
Koopman (Poly2+Cross+Trig+Bilin) 0.027 0.166 0.45 61 0.089 0.097 0.116
Koopman (Poly2+Cross+Bilin) 0.028 0.166 0.31 49 0.089 0.098 0.117
Koopman (MLP, 16) 0.028 0.167 205.3 16 0.078 0.087 0.141
Koopman (MLP, 24) 0.028 0.167 206.3 24 0.077 0.088 0.141
Koopman (Poly3+Cross+Trig) 0.029 0.170 0.326 49 0.094 0.101 0.138
Koopman (Poly2+Cross+Trig) 0.029 0.170 0.259 43 0.094 0.101 0.141
Koopman (Poly3+Cross) 0.029 0.170 0.231 37 0.094 0.101 0.141
Koopman (Poly2+Cross) 0.029 0.170 0.156 31 0.094 0.102 0.142
Baseline Models
Linear 0.033 0.182 0.010 0.140 0.149 0.144
Component-wise Linear 0.034 0.183 0.020 0.140 0.149 0.144
MLP 0.036 0.189 142.2 0.147 0.153 0.148
Integrator 0.083 0.289 0.178 0.173 0.236
TABLE I: Average prediction error on the validation dataset. The best performance is highlighted in bold.
Multi-layer Perceptron (MLP): We train MLP to learn the input and output of the transition function, as described in Sec. IV-C. Our network has two hidden layers of [32, 32] with a ReLU activation function.
Koopman: Our Koopman dynamics model has been tested with various lifting functions, including polynomial, cross, bilinear, and trigonometric terms. We also investigated MLP-based feature lifting functions, which simultaneously learns MLP feature layers and the transition matrix similar to the training process of the MLP model.
@@ V-B Long-term Forward Prediction
Figure 4: Long-term prediction error across the entire validation set. The bold line indicates the mean, and the shaded region shows the standard deviation.
Figure 5: Comparison of a long-term prediction of Koopman model against the baselines ($\Delta t=0.5$). The ground-truth trajectory is generated from a constant command not included in the training dataset. Koopman Forward Dynamics achieve two to four times less final step error compared to Linear and MLP baselines.
We train various Koopman dynamics models and compare them against several baselines, as summarized in Table I. The evaluation metrics include MSE, RMSE, training time, lifted dimension, and selected per-dimension RMSE for interpretability. Note that RMSE is computed in the 66D space for all the position and velocity values. Overall, these results demonstrate that Koopman lifting consistently outperforms baseline approaches in predictive accuracy. The full Koopman model with polynomial, cross, bilinear, and trigonometric observables attains the best RMSE score of $0.166$. Even the weakest Koopman variant with polynomial and cross terms achieves an RMSE of $0.170$, outperforming the most accurate baseline model, Linear, with an RMSE of $0.182$. While the Koopman model with learned MLP features yields the lowest RMSE along the X and Y directions, its inherent nonlinearity may complicate the MPC optimization process in the following section.
For the remainder of the experiments, we adopt the Koopman (Poly3+Cross+Trig) model, which achieves competitive performance while strictly adhering to the setup in Section III of $\psi(\mathrm{x}_{t},\mathrm{u}_{t})=[\phi(x_{t}),u_{t}]$; in particular, no bilinear state-control terms $\mathrm{x}_{i}\mathrm{u}_{j}$ are included in the lifting.
Although the performance gain reported in Table I may appear small, we want to note that they correspond to single-step predictions, and such errors accumulate rapidly in long-horizon rollouts. Fig. 4 presents long-term prediction errors across the validation set, where trajectories are predicted for 12 steps (6 seconds) using different models. While all models start with similar errors around 0.15, the Koopman model achieves a significantly lower long-term error of 0.37 after 12 steps, compared to 0.55 and 0.52 for the MLP and linear models, respectively. These results highlight the superior long-term prediction accuracy of the learned Koopman dynamics.
Fig. 5 illustrates a representative trajectory from the validation set, comparing predictions from Koopman, MLP, linear, and integrator models. While the baseline models deviate significantly after only a few steps, the Koopman model closely follows the ground-truth trajectory and achieves a final-step error that is two to four times lower than the others.
@@ V-C MPC-based Safe Navigation
In this section, we evaluate the performance of the proposed safe navigation framework (Sec. IV-D). We first evaluate path-tracking accuracy in obstacle-free environments and then examine navigation performance in corridor and maze environments.
V-C1 Pathtracking in obstacle free environments
Figure 6: Comparison of path-tracking trajectories (top), average distance from the reference path (bottom left), and solver time (bottom right).
We compare safe navigation with the Koopman model against two baselines: Linear and MLP. The Linear model is included as the best-performing baseline, while the MLP model is chosen for its ability to capture nonlinear dynamics. The three models are evaluated on two trajectories, a square and a circle, both involving multiple turns. Fig. 6 shows the reference and actual trajectories (top), along with the average distance from the reference path (bottom left) and the MPC solver time (bottom right).
In our experiments, we observe that navigation with the Koopman model achieves the best overall performance, as its trajectories almost overlap with the reference paths. In contrast, the Linear model shows noticeable deviations, particularly at sharp turns along the edges, highlighting its difficulty in accurately predicting turning behavior. The MLP-based model performs the worst, producing unstable and erratic trajectories. This is largely because incorporating an MLP as the dynamics model introduces highly nonlinear constraints into the MPC optimization, making it difficult for the solver to find a stable solution in real time. Quantitatively, our model achieves the lowest tracking errors of $0.039$ m and $0.041$ m in the square and circle tracking tasks, which correspond to 67.2% and 58.6% of the $0.058$ m and $0.070$ m errors of the second best model, Linear. These results demonstrate that the enhanced predictive accuracy of our learned Koopman forward dynamics directly translates into better navigation performance.
V-C2 Safe navigation in corridors and mazes
In this section, we compare the navigation performance of MPC with the learned Koopman dynamics against the Integrator and Linear models. The MLP model is omitted due to its poor tracking performance and high computational cost.
@@ V-C2 Safe navigation in corridors and mazes
We evaluate the effectiveness of the learned models in four navigation environments: i) Corridor 1, a wide corridor with a 70 cm gap; ii) Corridor 2, a narrow corridor with a 55 cm gap; iii) Maze 1, an artifact maze [25] with 0.9 m passages; and iv) Maze 2, the same maze with narrower passages of 0.8 m. For each environment, we randomize start and goal positions and sample five pairs. To isolate planning performance from perception, we assume obstacle positions and shapes are known. A high-level planner generates rough waypoints using the A* algorithm [46], which serve as reference trajectories for the MPC controller.
Corridor 1 Corridor 2 Maze 1 Maze 2
Time Succ. Time Succ. Time Succ. Time Succ.
Integrator 9.48 55 9.45 60 35.5 40 35.9 40
Linear 7.64 75 7.83 80 33.8 40 38.0 20
Koopman 8.63 75 8.88 80 32.7 80 29.7 100
TABLE II: Safe navigation results in four environments, where we compare the time (s) $\downarrow$ and success rate ($\%$) $\uparrow$.
@@
We measure the seconds taken to reach destination and the success rate of the navigation and report them in Table II. Here, the success is determined based on whether the robot was able to reach the destination in a desired time window, where we give $10$ seconds for corridors and $40$ seconds for maze environments. We did not terminate the episode upon minor collisions. However, when the robot fell down and was not be able to recover, we marked them as failure cases.
Overall, navigation with the Koopman model achieves the best performance, attaining the highest success rates across all four environments. In particular, it reaches success rates of 80% and 100% in Maze 1 and Maze 2, respectively, significantly outperforming the 40% and 20% achieved by the Linear model. This improvement arises because Koopman models are better at capturing turning motions compared to the baselines, and the maze environment requires frequent turning to complete the task. However, in terms of completion time in both Corridor 1 and Corridor 2, the Koopman models achieve the second-best performance, while the Linear model shows the fastest travel time. We observed that in some runs, collisions in the Linear model triggered aggressive behavior, allowing the robot to reach the target more quickly. However, this behavior is detrimental in the maze task, where successful navigation requires precise handling of complex turns.
Figure 8: Trajectory of safe navigation in Corridor 1. (a) The Koopman model successfully passes through the corridor without collision, while (b) the Linear baseline collides with the walls.
Fig. 8 illustrates a representative scenario in Corridor 1 by comparing the base trajectories of the Koopman and Linear models. The accurate predictive capability of the learned Koopman dynamics enables MPC-based navigation to follow the reference trajectory closely while avoiding collisions with nearby obstacles. In contrast, navigation with the Linear model overshoots at turns and results in severe collisions with the wall. These results highlight the potential of Koopman-based dynamics to improve safety in complex navigation tasks.